# 2Shanghai AI Lab, Shanghai, China

### v1

@@ 1 Introduction
Automatic and accurate delineation of the prostate plays an important role in assisting the diagnosis and treatment of prostate diseases. Despite that deep learning models have achieved remarkable performance on this task[16, 13], they often struggle to generalize well when confronted with gaps between training and testing data. To tackle this issue, Domain Adaptation (DA) methods emerge as a promising solution[5]. Unsupervised Domain Adaptation (UDA) has demonstrated considerable efficacy by leveraging knowledge from labeled source domain data to facilitate segmentation on unlabeled target domain[11, 29]. Moreover, given the constraints posed by privacy and security concerns, the unavailability of source domain necessitates extensive exploration of Source-Free Domain Adaptation (SFDA) techniques in medical image segmentation[18, 6, wu2023uplaggregate, 30]. Nonetheless, owing to the unforeseeable domain discrepancies, both UDA and SFDA face challenges in achieving satisfactory results.
@@ 3.1 Experimental Details
To further investigate each component’s contribution, we conducted ablation and sensitivity study on the first fold. The capacity $N_{tu}$ of the uncertainty candidate set during active sample selection is a hyper-parameter of our method. We set it to $M$, $2M$, $4M$, and $8M$ to investigate how it affects the performance, where $M=5\%$, in Fig.3(a). The results from one fold of cross-validation in both domains show that $4M$ is the best hyper-parameter to trade off performance and computational overhead. Further, in Fig.3(b), we also validate the effectiveness of our GAUA compared to other uncertainty estimation methods, including random, MC-dropout[10], Least Confidence(LC)[12] and highest entropy(Entropy)[23] followed typical practice of averaging the uncertainty across all the pixels to obtain image-level uncertainty. Our GAUA has achieved the highest performance, and all the methods’ performance have been buffed from TST. To demonstrate the necessity of utilizing the source model in the adaptation stage, we employed a few semi-supervised learning methods[7, 28], using 5% annotated data selected by our active sample selection technique for Semi-supervised Learning (SSL) in the HK domain under different training epochs. The results are presented in Fig.3(c). The performance of both stages of UGTST surpasses the existing SSL methods, demonstrating the priority and efficiency of DA in ASFDA. Next, we further validated the contribution of each component of our method in the domain BIDMC. The baseline involved using the source model’s predictions as pseudo-labels for adaptation. “Augmentation” means using the ensemble prediction as the pseudo-label to apply self-training process without annotation. When not using TST, we directly merge active samples with labels and unlabeled samples with pseudo labels for self-training. Experimental results shown in Table 2 show marked performance improvements by each component of UGTST, further confirming the effectiveness of our approach.

### v2

@@ 1 Introduction
Automatic and accurate delineation of the prostate plays an important role in assisting the diagnosis and treatment of prostate diseases. Despite that deep learning models have achieved remarkable performance on this task[15, 12], they often struggle to generalize well when confronted with gaps between training and testing data[18]. To tackle this issue, Domain Adaptation (DA) methods emerge as a promising solution[5]. Unsupervised Domain Adaptation (UDA) has demonstrated considerable efficacy by leveraging knowledge from labeled source domain data to facilitate segmentation on unlabeled target domain[27, 30]. Moreover, given the constraints posed by privacy and security concerns, the unavailability of source domain necessitates extensive exploration of Source-Free Domain Adaptation (SFDA) techniques in medical image segmentation[17, 6, wu2023uplaggregate, 31]. Nonetheless, owing to the unforeseeable domain discrepancies, both UDA and SFDA face challenges in achieving satisfactory results.
@@ 3.1 Experimental Details
To further investigate each component’s contribution, we conducted ablation and sensitivity study on the first fold. The capacity $N_{tu}$ of the uncertainty candidate set during active sample selection is a hyper-parameter of our method. We set it to $M$, $2M$, $4M$, and $8M$ to investigate how it affects the performance, where $M=5\%$, in Fig.3(a). The results from one fold of cross-validation in both domains show that $4M$ is the best hyper-parameter to trade off performance and computational overhead. Further, in Fig.3(b), we also validate the effectiveness of our GAUA compared to other uncertainty estimation methods, including random, MC-dropout[10], Least Confidence(LC)[11] and highest entropy(Entropy)[23] followed typical practice of averaging the uncertainty across all the pixels to obtain image-level uncertainty. Our GAUA has achieved the highest performance, and all the methods’ performance have been buffed from TST. To demonstrate the necessity of utilizing the source model in the adaptation stage, we employed a few semi-supervised learning methods[7, 29], using 5% annotated data selected by our active sample selection technique for Semi-supervised Learning (SSL) in the HK domain under different training epochs. The results are presented in Fig.3(c). The performance of both stages of UGTST surpasses the existing SSL methods, demonstrating the priority and efficiency of DA in ASFDA.
Next, we further validated the contribution of each component of our method in the domain BIDMC. The baseline involved using the source model’s predictions as pseudo-labels for adaptation. “Augmentation” means using the ensemble prediction as the pseudo-label to apply self-training process without annotation. When not using TST, we directly merge active samples with labels and unlabeled samples with pseudo labels for self-training. Experimental results shown in Table 2 show marked performance improvements by each component of UGTST, further confirming the effectiveness of our approach.